OpenStudy (anonymous):

y=sqrt(x), find dy/dt when x=4, given dx/dt=3.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=ywEqxpgxark

OpenStudy (jdoe0001):

keep in mind that \(\large \bf \sqrt{x}\implies x^{\frac{1}{2}}\)

OpenStudy (anonymous):

Can someone explain this to me step by step?

OpenStudy (anonymous):

utube linlk

OpenStudy (anonymous):

link*

OpenStudy (anonymous):

its similar

zepdrix (zepdrix):

Ok let's start with this: \[\Large \frac{dy}{dx}\quad=\quad \frac{\left(dy/dt\right)}{\left(dx/dt\right)}\]

OpenStudy (anonymous):

@UchihaMadara I see.

OpenStudy (anonymous):

could i get a medal

zepdrix (zepdrix):

Given a function y, we should be able to find dy/dx without too much trouble. Then plug in dy/dx evaluated at x=4, and plug in dx/dt. Andddd move some stuff around to solve for dy/dt.

OpenStudy (anonymous):

could i get a medal

OpenStudy (anonymous):

\[y=x ^\frac{ 1 }{2 },\frac{ dy }{dt }=\frac{ 1 }{2 }x ^{\frac{ -1 }{ 2 }}\frac{ dx }{ dt }\] \[when x=4,\frac{ dx }{ dt }=3\] \[\frac{ dy }{ dt }=\frac{ 1 }{2 }\left( 4 \right)^{\frac{ -1 }{2 }}*3=\frac{ 3 }{2*2 }=\frac{ 3 }{4 }\]